Given the equation p = q^2 + 5q +2. Find the values of q corresponding to
p = -2.

Answer by nerdybill(7384)

Given the equation p = q^2 + 5q +2. Find the values of q corresponding to
p = -2.
.
Start with:
p = q^2 + 5q +2
Stuff in out value of p:
-2 = q^2 + 5q +2
Add two to both sides:
0 = q^2 + 5q + 4
Factoring the right side:
0 = (q+4)(q+1)
.
q = {-4, -1}